### v3

@@ 1. Introduction
The integer $\langle u,u^{\prime}\rangle_{w}$ only depends on $u$ and $u^{\prime}$, not on the choice of seed $({\bf x}_{w},\widetilde{B}_{w})$.
From the proof of Theorem 4.8, we can see that the integer $\langle u,u^{\prime}\rangle_{w}$ can be interpreted as the evaluation of the positive function $u$ (on a cluster $\mathscr{A}$-variety in the sense of [FG-2009]) at a tropical point (associated to the same cluster $\mathscr{A}$-variety) determined by $u^{\prime}$. Different choices of seeds provide different charts to calculate this evaluation.
By Theorem 1.1, we get an invariant $\langle u,u^{\prime}\rangle:=\langle u,u^{\prime}\rangle_{w}$ for any pair $(u,u^{\prime})$ of good elements. We call $\langle u,u^{\prime}\rangle$ the tropical invariant between $u$ and $u^{\prime}$. We remark that the tropical invariant $\langle u,u^{\prime}\rangle=\langle u,u^{\prime}\rangle_{w}$ is an analog of the invariant $\Lambda(M,N)$ in the monoidal categorification of cluster algebras, because both $\langle u,u^{\prime}\rangle$ and $\Lambda(M,N)$ are related to the quantization matrices appearing in compatible pairs that are used to define quantum cluster algebras in the sense of [bz-2005], cf. Remark 4.14.
The $F$-invariant $(u\mid\mid u^{\prime})_{F}$ is defined to be the symmetrized sum of the tropical invariants.
(1.2) $(u\mid\mid u^{\prime})_{F}:=\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle=\langle u,u^{\prime}\rangle_{w}+\langle u^{\prime},u\rangle_{w}.$
We remark that the $F$-invariant defined above is an analog of the $E$-invariant $E^{\rm sym}(\mathcal{M},\mathcal{N})$ in the additive categorification and $\mathfrak{d}$-invariant $\mathfrak{d}(M,N)$ in the monoidal categorification. Strictly speaking, the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ is analogous to $2\mathfrak{d}(M,N)$, because in (1.2) we did not divide out ``$2"$ from the symmetrized sum $\langle u,u^{\prime}\rangle+\langle u^{\prime},u\rangle$.
We prove that the $F$-invariant admits the similar properties with the $E$-invarant and $\mathfrak{d}$-invariant.
Proposition 1.2 (Proposition 4.11).
The following statements hold.
@@
    For any pair $(u,u^{\prime})$ of good elements, we have $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
@@
    If $u$ is a cluster monomial, then $(u\mid\mid u)_{F}=0$.
We remark that the class of good elements satisfying $(u\mid\mid u)_{F}=0$ is analogous to the class of decorated representations satisfying $E^{\rm sym}(\mathcal{M},\mathcal{M})=0$ and to the class of finite-dimension graded modules satisfying $\mathfrak{d}(M,M)=0$.
As mentioned in [kkop-2020, Pages 1040–1041], the integer-valued invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ provide important information in the representation theory of quiver Hecke algebra. However, in general, computing those values is quite difficult. We hope our tropical invariant $\langle u,u^{\prime}\rangle$ and $F$-invariant $(u\mid\mid u^{\prime})_{F}$ can shed some light on the calculation of the integer-valued invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ in the monoidal categorification of cluster algebras.
Since the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ always takes non-negative values, it is natural to ask when $(u\mid\mid u^{\prime})_{F}=0$ holds.
@@
We use $m$ to denote the number of cluster variables (frozen and unfrozen) in a cluster and use $n$ to denote the number of unfrozen cluster variables in a cluster. By Definition 3.6, each good element $u$ corresponds to a tropical point $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ (associated to a cluster $\mathscr{X}$-variety in the sense of [FG-2009]). Two tropical points $[{\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and $[{\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ are said to be sign-coherent if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
@@
Let $u$ and $u^{\prime}$ be two good elements in an upper cluster algebra $\mathcal{U}$. If $(u\mid\mid u^{\prime})_{F}=0$, then the tropical points correspond to $u$ and $u^{\prime}$ are sign-coherent and the product $uu^{\prime}$ is still a good element in $\mathcal{U}$.
We also have a discussion on the link between the compatible Poisson structure of (upper) cluster algebras and the $F$-invariant.
@@
The following statements hold.
@@
    If a good element $u$ and an unfrozen cluster variable $x$ are log-canonical with respect to the given compatible Poisson bracket, then $(x\mid\mid u)_{F}=0$.
@@
Dominant sets and oriented exchange graphs. Two seeds of a cluster algebra are said to be equivalent, if they are the same up to a relabeling. The exchange graph $\mathcal{H}=\mathcal{H}(\mathcal{A})$ of a cluster algebra $\mathcal{A}$ is the graph whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose edges correspond to seed mutations. Thanks to the sign-coherence of $C$-matrices [GHKK18], each mutation has two states: green mutation or red mutation. Fix an initial seed $({\bf x}_{t_{0}},\widetilde{B}_{t_{0}})$, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ is the quiver whose underlying graph is the exchange graph $\mathcal{H}$ of $\mathcal{A}$ and whose orientation is induced by the green mutations.
@@
$\tau$-tilting theory is introduced by Adachi, Iyama and Reiten [air_2014] as a far reaching generalization of cluster-tilting theory (one of the first major developments in representation theory discovered thanks to the relation with cluster algebras). Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $C$-matrices, $G$-matrices, (oriented) exchange graphs, etc., were generalized to $\tau$-tilting theory.
@@
The seeds in cluster algebras correspond to (basic) $\tau$-tilting pairs in $\tau$-tilting theory, which are in bijection with the functorially finite torsion classes in the module categories. The inclusion of subcategories induces a natural partial order on the set of functorially finite torsion classes and thus a partial order on the set of $\tau$-tilting pairs (up to isomorphisms).
@@
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as an replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory and as an replacement of inversion sets in the study of right weak order on Weyl groups. With the help of the dominant sets, we prove the following results on oriented exchange graphs.
Theorem 1.6 (Theorems 5.7, 5.12, Corollary 5.13).
@@
Remark 1.7.
When I was a PhD student (2014-2019), I was interested in the above problems. But at that time, I was not able to prove them, because I did not know what is the right replacements of torsion classes in cluster algebras.
@@
Finally, we would like to point out that by ignoring initial cluster variables, the dominant sets in cluster algebras correspond to the sets of indecomposable $\tau$-rigid modules contained in the functorially finite torsion classes (see Proposition 5.21). In this sense, we view the dominant sets as an replacement of torsion classes in $\tau$-titling theory and please see Section 5.4 for details. By ignoring initial cluster variables, the dominant sets also correspond to the inversion sets in the study of right weak order on Weyl groups (see Proposition 5.30).
Organization of the paper. In Section 2, we recall some well-known concepts in cluster algebras, including $Y$-patterns, cluster patterns, compatible pairs, and (upper) cluster algebras. In Section 3, we recall the definition of tropical points for $Y$-patterns (equivalently, cluster $\mathscr{X}$-varieties) and for cluster patterns (equivalently, cluster $\mathscr{A}$-varieties), and we also recall the (compatibly) pointed elements introduced by Qin [Qin_2017, qin_2019]. We introduce the notion of good elements in upper cluster algebras. The $F$-invariant will be defined for a pair of good elements.
In Section 4.1, we give two maps between the set of tropical points. Let $(\mathcal{S}_{X},\mathcal{S}_{Y})$ be a Langlands dual pair in the sense of Definition 2.6. The first map is from the set of tropical points associated to the cluster pattern $\mathcal{S}_{X}$ to the set of tropical points associated to the $Y$-pattern $\mathcal{S}_{Y}$ and the second map is from the set of tropical points associated to the $Y$-pattern $\mathcal{S}_{Y}$ to the set of tropical points associated to the cluster pattern $\mathcal{S}_{X}$. We need to emphasise that the second map will play an important role in the proof Theorem 4.8 which says that the pairing $\langle u,u^{\prime}\rangle_{w}$ is independent of the choice of seed $({\bf x}_{w},\widetilde{B}_{w})$.
In Section 4.2, we give the definition of the tropical invariant and $F$-invariant. The main results in this subsection are Theorem 4.8, Theorem 4.22 and Theorem 4.27. In Section 4.3, we discuss the link between the compatible Poisson structure of (upper) cluster algebras and the $F$-invariant and the main result in this subsection is Theorem 4.29.

### v4

@@ 1.2. Results (I): Tropical invariant and $F$-invariant
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. Let
$u={\bf x}_{w}^{{\bf g}_{u}^{w}}F_{u}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w}),\;\;\;u^{\prime}={\bf x}_{w}^{{\bf g}_{u^{\prime}}^{w}}F_{u^{\prime}}^{w}(\hat{y}_{1;w},\ldots,\hat{y}_{n;w})$
be the canonical expressions of $u$ and $u^{\prime}$ with respect to any vertex $w\in\mathbb{T}_{n}$. Then we have
@@
    $\langle u,u^{\prime}\rangle_{\rm trop}=({\bf g}_{u}^{w})^{T}\Lambda_{w}{\bf g}_{u^{\prime}}^{w}+F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$, where $(S\mid{\bf 0})$ is as in (1.1).
@@
    $(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}]$. In particular, $(u\mid\mid u^{\prime})_{F}\in\mathbb{Z}_{\geq 0}$.
  • (iii)
    $(u\mid\mid u)_{F}=0$ whenever $u$ is a cluster monomial.
Remark 1.4.
From the formula in (i), we see that the tropical invariant heavily depends on the Poisson coefficient matrix $\Lambda_{w}$. Since we define $F$-invariant by taking the symmetrized sum of tropical invariants, a prior, the $F$-invariant should also heavily depend on the Poisson coefficient matrix $\Lambda_{w}$. However, if we look at the formula in (ii), this formula eliminates the dependence of $(u\mid\mid u^{\prime})_{F}$ on the Poisson coefficient matrix and it only depends on the notion of $g$-vectors and $F$-polynomials. From this viewpoint, $F$-invariant can be defined for any skew-symmetrizable (upper) cluster algebra with trivial coefficients if we restrict to cluster monomials (see Proposition 4.30). The advantange to work on $\mathsf{\Lambda}$-(upper) cluster algebras is that we can define $F$-invariant for any pair of good elements and go beyond the class of cluster monomials.
Since the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ can be written as
$(u\mid\mid u^{\prime})_{F}=F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]+F_{u^{\prime}}^{w}[(S\mid{\bf 0}){\bf g}_{u}^{w}],$
we call the non-negative integer $F_{u}^{w}[(S\mid{\bf 0}){\bf g}_{u^{\prime}}^{w}]$ the partial $F$-invariant of the pair $(u,u^{\prime})$ at vertex $w\in\mathbb{T}_{n}$.
As the $F$-invariant $(u\mid\mid u^{\prime})_{F}$ always takes non-negative values, it is natural to ask when $(u\mid\mid u^{\prime})_{F}=0$ holds.
@@
Recall that $m$ is the number of cluster variables (frozen and unfrozen) in a cluster and $n$ is the number of unfrozen cluster variables in a cluster. By Definition 3.6, each good element $u$ corresponds to a tropical point ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ (associated to a $Y$-pattern). Two tropical points ${\bf g}]=\{{\bf g}^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ and ${\bf g}^{\prime}]=\{({\bf g}^{\prime})^{t}\in\mathbb{Z}^{m}\mid t\in\mathbb{T}_{n}\}$ are said to be sign-coherent if for any vertex $t\in\mathbb{T}_{n}$ and any $k\in[1,n]$, we have $g_{k}^{t}\cdot(g_{k}^{\prime})^{t}\geq 0$, where $g_{k}^{t}$ and $(g_{k}^{\prime})^{t}$ are the $k$th components of ${\bf g}^{t}$ and $({\bf g}^{\prime})^{t}$.
@@
Let $u$ and $u^{\prime}$ be two good elements in a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$. If $(u\mid\mid u^{\prime})_{F}=0$, then the tropical points correspond to $u$ and $u^{\prime}$ are sign-coherent and the product $uu^{\prime}$ is still a good element in $\mathcal{U}$.
Since we work on a $\mathsf{\Lambda}$-upper cluster algebra $\mathcal{U}$, it has a natural compatible Poisson structure. By definition, if two cluster variables are in the same cluster, then they are log-canonical with respect to the compatible Poisson structure. We prove that the converse statement is also true.
@@
Let $\mathcal{U}$ be a $\mathsf{\Lambda}$-upper cluster algebra. Then the following statements hold.
@@
    Let $u\in\mathcal{U}$ be a good element and $x\in\mathcal{U}$ an unfrozen cluster variable. If $u$ and $x$ are log-canonical, then $(x\mid\mid u)_{F}=0$.
@@
1.3. Results (II): Comparison with some categorical invariants
As we have seen, the $F$-invariant shares the similar properties with the $E$-invariant in the additive cluster categorification and the $\mathfrak{d}$-invariant in the monoidal cluster categorification. It is natural to compare them.
We first focus on the monoidal cluster categorification [HL_2010, HL_2013]. A finite length, $\mathbf{k}$-linear, monoidal abelian category $\mathcal{C}$ is a monoidal categorification of a cluster algebra $\mathcal{A}^{+}$ (with non invertible frozen variables), if there exists a $\mathbb{Z}$-algebra isomorphism
$\varphi\colon\mathcal{K}_{0}(\mathcal{C})\rightarrow\mathcal{A}^{+}$
from the Grothendieck ring $\mathcal{K}_{0}(\mathcal{C})$ to the cluster algebra $\mathcal{A}^{+}$ such that the cluster monomials of $\mathcal{A}^{+}$ correspond to a subset of the simple objects (up to isomorphism) in $\mathcal{C}$. Such simple objects are said to be reachable. We denote by $(\mathcal{A}^{+},\mathcal{C},\varphi)$ for this monoidal categorification.
In order to unify the definition of $\Lambda$-invariant and $\mathfrak{d}$-invariant in different settings and unify the proofs for our results, we introduce the notion of $\mathsf{\Lambda}$-structure on a monoidal (abelian) category, which extracts the key properties of $\Lambda$-invariant appearing in the representation theory of quiver Hecke algebras [kkko-2018] and that of quantum affine algebras [kkop-2020].
Roughly speaking, a $\mathsf{\Lambda}$-structure on a monoidal category $\mathcal{C}$ is a pair $\mathsf{\Lambda}=(\Omega,\Lambda)$, where
  • $\Omega$ is a class of non-zero objects in $\mathcal{C}$ containing all the simple objects and closed under isomorphism, taking tensor products and non-zero subquotients;
  • ${\Lambda}:\Omega\times\Omega\rightarrow\mathbb{Z}$ is a map satisfying certain axiomatic conditions (see Definition 5.5).
If $\mathsf{\Lambda}=(\Omega,\Lambda)$ is a $\mathsf{\Lambda}$-structure on a monoidal category $\mathcal{C}$, we simply say that $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ is a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. For $M,N\in\Omega$, the numbers $\Lambda(M,N)$ and
$\mathfrak{d}(M,N)\coloneqq\frac{\Lambda(M,N)+\Lambda(N,M)}{2}$
are respectively called the $\Lambda$}-invariant and $\mathfrak{d}$-invariant of the pair $(M,N)$.
Definition 1.8 ($\mathsf{\Lambda}$-monoidal categorification [^1] We borrow this teminology from [kkop-2020, Definition 6.7].).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi)$ be a monoidal categorification and $\Lambda\colon\Omega\times\Omega\rightarrow\mathbb{Z}$ a $\mathsf{\Lambda}$-structure on $\mathcal{C}$. We say that $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ is a $\mathsf{\Lambda}$}-monoidal categorification if the following conditions hold.
  • (c1)
    Let $M,N\in\mathcal{C}$ be simple objects such that at least one of them is real. Then the head $M\nabla N$ of $M\otimes N$ is simple.
  • (c2)
    The $\mathsf{\Lambda}$-structure and cluster structure on $\mathcal{C}$ are compatible in the sense that there exists a monoidal cluster $\mathcal{M}_{w}=(M_{1;w},\ldots,M_{m;w})$ in $\mathcal{C}$ such that $(\widetilde{B}_{w},\Lambda_{w})$ forms a compatible pair, where $\widetilde{B}_{w}$ comes from the cluster structure and $\Lambda_{w}=(\lambda_{ij;w})$ with $\lambda_{ij;w}=\Lambda(M_{i;w},M_{j;w})$ comes from the $\mathsf{\Lambda}$-structure.
The examples of $\mathsf{\Lambda}$-monoidal categorifications contain those monoidal categorifications constructed by Kang, Kashiwara, Kim, and Oh [kkko-2018] using monoidal subcategories of representations of (symmetric) quiver Hecke algebras and by Kashiwara, Kim, Oh, and Park [kkop-2024] using monoidal subcategories of representations of quantum affine algebras.
Thanks to Proposition 5.13, the cluster algebra $\mathcal{A}^{+}$ above, together with the family of skew-symmetric matrices $\{\Lambda_{t}\mid t\in\mathbb{T}_{n}\}$ forms a $\mathsf{\Lambda}$-cluster algebra. Thus the tropical invariant $\langle u,u^{\prime}\rangle_{\rm trop}$ is defined for any pair of good elements in $\mathcal{A}^{+}$.
Theorem 1.9 (Theorem 5.16).
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. Let $M,N$ be two simple objects in $\mathcal{C}$ such that $\varphi(M)$ and $\varphi(N)$ are good elements in $\mathcal{A}^{+}$. If $M$ or $N$ is reachable, then we have
$\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}\;\;\;\text{and}\;\;\;2\mathfrak{d}(M,N)=(\varphi(M)\mid\mid\varphi(N))_{F}.$
Since the reachable simple objects correspond to cluster monomials, which are always good elements in $\mathcal{A}^{+}$, we have the following consequence.
Corollary 1.10.
Let $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ be a $\mathsf{\Lambda}$-monoidal categorification. If $M$ and $N$ are reachable simple objects, then we have
$\Lambda(M,N)=\langle\varphi(M),\varphi(N)\rangle_{\rm trop}\;\;\;\text{and}\;\;\;2\mathfrak{d}(M,N)=(\varphi(M)\mid\mid\varphi(N))_{F}.$
As mentioned in [kkop-2020, Pages 1040–1041], the integer-valued invariants $\Lambda(M,N)$ and $\mathfrak{d}(M,N)$ provide important information in the representation theory of quiver Hecke algebra. However, in general, computing those values is quite difficult. Thanks to Theorem 1.9 and Theorem 1.3, we obtain explicit formulas expressing the $\Lambda$-invariant and $\mathfrak{d}$-invariant via $g$-vectors and $F$-polynomials, which can be easily implemented by a computer program.
Now, let us turn to the additive cluster categorification [DWZ08, DWZ10].
Theorem 1.11 (Theorem 6.11).
Let $(Q,W)$ be a quiver with non-degenerate potential and $\mathcal{A}_{Q}$ the corresponding cluster algebra with trivial coefficients. Let $u,u^{\prime}$ be two cluster monomials of $\mathcal{A}_{Q}$ and let $\mathcal{M}_{u},\mathcal{M}_{u^{\prime}}$ be the negative-reachable decorated representations of $(Q,W)$ corresponding to $u$ and $u^{\prime}$. Then
$(u\mid\mid u^{\prime})_{F}=E^{\rm sym}(\mathcal{M}_{u},\mathcal{M}_{u^{\prime}}).$
Notice that in the above theorem we take $S=I_{n}$ to be the fixed skew-symmetrizer for the exchange matrices of $\mathcal{A}_{Q}$, when we consider the $F$-invariant for cluster monomials in $\mathcal{A}_{Q}$.
1.4. Results (III): Dominant sets and oriented exchange graphs
@@
Two seeds of a cluster algebra are equivalent, if they are the same up to relabeling. The exchange graph $\mathcal{H}=\mathcal{H}(\mathcal{A})$ of a cluster algebra $\mathcal{A}$ is the graph whose vertices correspond to the seeds (up to seed equivalence) of $\mathcal{A}$ and whose edges correspond to seed mutations. Thanks to the sign-coherence of $C$-matrices [GHKK18], each mutation has two states: green mutation or red mutation. Fix an initial seed at vertex $t_{0}\in\mathbb{T}_{n}$, the oriented exchange graph $\overrightarrow{\mathcal{H}}^{t_{0}}$ of $\mathcal{A}$ is the quiver whose underlying graph is the exchange graph $\mathcal{H}$ of $\mathcal{A}$ and whose orientation is induced by the green mutations.
@@
$\tau$-tilting theory is introduced by Adachi, Iyama and Reiten [air_2014], which completes (classical) tilting theory from the viewpoint of mutation. Various of the fundamental concepts in cluster algebras, such as clusters/seeds, mutations, $C$-matrices, $G$-matrices, (oriented) exchange graphs, etc., were generalized to $\tau$-tilting theory.
@@
The seeds in cluster algebras correspond to (basic) $\tau$-tilting pairs in $\tau$-tilting theory, which are in bijection with the functorially finite torsion classes in the module categories. The inclusion of subcategories induces a natural partial order on the set of functorially finite torsion classes and thus a partial order on the set of $\tau$-tilting pairs (up to isomorphism).
@@
Inspired by $F$-invariant, we introduce the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory. With the help of the dominant sets, we prove the following results on oriented exchange graphs.
Theorem 1.12 (Theorems 7.8, 7.13, Corollary 7.14).
@@
We would like to point out that by ignoring initial cluster variables, the dominant sets in cluster algebras correspond to the sets of indecomposable $\tau$-rigid modules contained in the functorially finite torsion classes (see Proposition 7.23). In this sense, we view the dominant sets as a replacement of torsion classes in $\tau$-titling theory.
1.5. Conclusion and future work
In this paper, we introduce the tropical invariant and $F$-invariant for ($\mathsf{\Lambda}$)-cluster algebras with a focus on $F$-invariant. Up to a constant rescaling factor, we prove that the following three invariants coincide on cluster monomials: (i) the $\mathfrak{d}$-invariant in $\mathsf{\Lambda}$-monoidal cluster categorification; (ii) the $F$-invariant in cluster algebras; (iii) the $E$-invariant in additive cluster categorification. Notice that these three invariants are symmetric, e.g., $(u\mid\mid u^{\prime})_{F}=(u^{\prime}\mid\mid u)_{F}$. We prove the oriented exchange graphs of cluster algebras are acyclic by introducing the dominant sets for seeds of cluster algebras as a replacement of torsion classes for $\tau$-tilting pairs in $\tau$-tilting theory.
Now let us consider those $\mathsf{\Lambda}$-monoidal categorifications $(\mathcal{A}^{+},\mathcal{C},\varphi,\Omega,\Lambda)$ constructed by Kashiwara, Kim, Oh and Park [kkop-2024] using the representations of quantum affine algebras. For a pair $(M,N)$ of simple objects in $\mathcal{C}$, it is known [kkop-2020, Proposition 3.16] that the $\mathfrak{d}$-invariant can be written as
$\mathfrak{d}(M,N)=\mathfrak{o}(M,N)+\mathfrak{o}(N,M),$
where $\mathfrak{o}(M,N)\in\mathbb{Z}_{\geq 0}$ is the pole order of the normalized $R$-matrix $R^{\mathrm{norm}}_{M,N_{z}}$ at $z=1$. We may call $\mathfrak{o}(M,N)$ the partial $\mathfrak{d}$-invariant of the pair $(M,N)$. It is an important quantity in itself, particularly as a key ingredient in the construction of the generalized quantum affine Schur–Weyl duality functors due to [KKK-2018R].
In an ongoing work [^2] This work will be an extended version of the note [cao_2025]. with Ryo Fujita and Kota Murakami, we show that the following three (non-symmetric) partial invariants coincide on cluster monomials with a good choice of the reference seed: (i) the partial $\mathfrak{d}$-invariant in $\mathsf{\Lambda}$-monoidal cluster categorification using representations of quantum affine algebras; (ii) the partial $F$-invariant in cluster algebras; (iii) the partial $E$-invariant in additive cluster categorification. Theorem 1.9 will play an important role in the proof of the coincidence of the partial $F$-invariant and partial $\mathfrak{d}$-invariant. Such results will be used to prove the conjectural unified denominator formula [Fujita-Oh-2021, Conjecture 6.7] (equivalently, [Fujita-Murakami-2023, Conjecture 5.17]) for the normalized $R$-matrices between the KR-modules.
1.6. Organization
In Section 2, we recall some well-known concepts in cluster algebras, including $Y$-patterns, cluster patterns, compatible pairs, and $\mathsf{\Lambda}$-(upper) cluster algebras.
In Section 3, we recall the definition of tropical points for $Y$-patterns and cluster patterns, and compatibly pointed elements in upper cluster algebras. We give the definition of good elements for upper cluster algebras of full rank.
In Section 4, for each good element in a $\mathsf{\Lambda}$-upper cluster algebra, we introduce a semifield homomorphism (see Proposition 4.4) and we define the tropical invariant and $F$-invariant using such semifield homomorphisms. The main results in this section are Theorem 4.10, Theorem 4.19 , Theorem 4.24, Theorem 4.26. Notice that $F$-invariant can be defined for any skew-symmetrizable (upper) cluster algebra with trivial coefficients, when we only consider cluster monomials (see Proposition 4.30).
In Section 5 and Section 6, we compare the tropical invariant and $F$-invariant with some known invariants in additive/monoidal cluster categorification. When restricting to cluster monomials, we prove that the tropical invariant and $F$-invariant respectively coincide with the $\Lambda$-invariant and twice $\mathfrak{d}$-invariant in $\mathsf{\Lambda}$-monoidal categorification (see Theorem 5.16) and we prove that the $F$-invariant coincides with the $E$-invariant in additive cluster categorification (see Theorem 6.11).